# Leveraging Domain Knowledge for Efficient Reward Modelling in RLHF: A Case-Study in E-Commerce Opinion Summarization

### v1

@@ 6 Analysis
We perform a two-fold analysis: (a) First, we see the domain knowledge feature influences for $\varphi_{op}$, (b) Second, we see how the ground truth summary and summary from trained models fare on the domain knowledge features. This two-fold analysis helps us understand: (a) which features influence the latent reward model within humans [^6] Note that the trained $\varphi_{op}$ represents latent human reward model. the most, and (b) how the ground truth summary and summary from trained models fare on these influential features. Performing good on influential features would mean the summary aligns well with the latent reward model within humans.
@@ 6.1 Analysis of $\varphi_{op}$
$\varphi_{op}$ model has been trained on a set of features specified by domain experts. We analyze the relative influence of each feature on the final score assigned by $\varphi_{op}$. Doing this helps us understand an approximate importance [^7] We call this approximate importance as influence of a features on the output is not necessarily its importance. of each of these features. We do this by varying each feature by $\delta\>(=0.1)$ while keeping the other features constant, over multiple possible values of all features (Equation 6.1).
@@
Figure 7 highlights the relative influence of all the features. We see that hallucination is most influential. This aligns with what our human preference annotators report – hallucination in a summary is the primary cause of rejection. We see that, in majority cases, hallucinations are within the opinions in the summary. We see that this is also reflected in Figure 7opinion-faithfulness has significant influence. We also see that annotators prefer summaries which have more specifics about a product, i.e. they include more aspects – aspect-coverage has significant influence.
@@ 6.2 Analysis of Summaries
We analyze the top-$3$ performing models (in human and Gpt-$4$ evaluations) for the following features: opinion-coverage, opinion-faithfulness, hallucination and relevance. We show the analysis only for Amazon benchmark in the main body of the paper, we include rest in Appendix H. Figure 6 shows the performance on these features. We see that Inductive-Bias model fares much better than the competitors on hallucination (the most influential metric). For relevance, aspect-coverage and opinion-faithfulness, our model is fairly better than the other models.
@@
This shows that our technique helps Inductive-Bias model perform well on features that influence the latent reward model within humans for opinion summarization. This means that our technique helps Inductive-Bias model achieve a significant alignment with the latent reward model. This conclusion verifies our hypothesis (in the domain of opinion summarization) – An inductive bias infused reward model ($\varphi_{op}$) can help achieve alignment with latent reward model of humans for a task, with modest human preference annotations.
@@ 7 Summary, Conclusion and Future Work
In this work, we propose a novel Reward Modelling technique via Domain Knowledge Infusion. We verify our approach for E-Commerce Opinion Summarization, where we achieve state-of-the-art, while significantly reducing the amount of human preference annotations required (just $940$ samples). In addition to advancing SOTA and reducing preference annotations, our technique provides another two-fold benefits: (a) No Alignment Tax, and (b) Interpretability. Due to the interpretable nature, we find that our model does achieve alignment with human goals for Opinion Summarization through analysis. From the results and analysis, we conclude that Domain Knowledge Infusion into Reward Modelling is a viable solution to reduce human preference annotations for downstream tasks. In future, we would verify this for other domains.
@@ 8 Ethical Considerations
We contribute two datasets in our work: PromptOpinSumm, OpinPref. These datasets are generated using an open-source model Mistral-$7$B Jiang et al. (2023). We would release the datasets to further research in Opinion Summarization. For the OpinPref, we have seen that it does not contain any harmful content, such as social biases, stereotypes, etc. However, we have seen that it contains products of explicit nature (sexual products). For the PromptOpinSumm dataset, to the best of our knowledge, there is no presence of harmful content, such as social biases, stereotypes etc. We urge the research community to use the datasets with caution and check for potential harmfulness, based on their use-cases.
@@ 9 Limitations
A limitation of our work is we have tested our approach for one domain – Opinion Summarization. However, we do not believe that this weakens our argument, as we have exhaustively shown that our approach not only advances SOTA, but also interpretably achieves alignment with humans. Future work in other domains would help in verifying this claim for other domains. Another limitation is: we train our RLHF models using the limited trajectory trick. However, this limitation is imposed by the available compute resources. With larger compute resources, this study can be extended to extensive exploration during RLHF training.
Appendix A Features for Reward Modelling
We use $7$ domain specific features for the reward model $\varphi_{op}$. For each feature we prompt Mistral-$7$B to generate a score within 00 and $5$. For each feature, 00 means the model is doing bad on the feature, and $5$ means the model is doing good on the feature. We define all the features below:
@@ Appendix A Features for Reward Modelling
conciseness: This feature considers the conciseness and completeness of the opinion summary. The feature assumes a value $5$ if the summary is concise and complete – not one phrase/sentence can be dropped off. It assumes a value 00 if the summary is totally incomplete, or very verbose.
@@
Appendix B Example Summaries in Existing Benchmarks
We highlight the shortcomings of the benchmark datasets here.
@@ Appendix B Example Summaries in Existing Benchmarks
Amazon: Bražinskas et al. (2020) designed the test-set in such a way that the summary has to read like a review, for instance, summary would contain ‘I think the quality has come down over the years.’, instead of ‘Users think that quality has come down over years’. Due to this writing style, the summaries read like reviews and are often in first person – high overlap would not necessarily mean a better summary, it would rather mean a better review.
@@
Appendix C Implementation Details
@@
Appendix D Generated Summary Lengths
We analyze the generation lengths of the models, and the ground truth summary. Table 6 lists the summary lengths.
Model Amazon Oposum Flipkart
@@ Appendix D Generated Summary Lengths
NaiveMean $79.93$ $99.76$ $71.59$
Synth-Feedback $87.59$ $115.00$ $84.18$
@@
Inductive-Bias $73.53$ $68.63$ $68.63$
Supervised $81.84$ $126.00$ $77.80$
Table 6: Generation Length Statistics: number of words in summaries. We use NLTK to split into tokens.
@@
Table 7: Example summaries of generated dataset
Appendix E Details of the Generated Dataset
We include an example from PromptOpinSumm dataset (Table 7). We show one sample from Good-Sum, SBad-Sum and VBad-Sum quality each. We do not include the reviews for conciseness. However, we incluce the salient aspects of the reviews. The reviews talk about the following things:
@@ Appendix E Details of the Generated Dataset
Good-Sum manages to discuss all of these things – it was able to detect that the cover does not fit specific models (highlighted in blue). It also detects that a few users like the “personal touch addition” factor of the cover.
VBad-Sum is also a good representation of a bad summary – it totally ignores the “positive” aspects of the product and presents the “negative” aspects only. It fails at Aspect Coverage, Opinion Faithfulness and Opinion Coverage.
@@
Appendix F Statistics of the OpinPref dataset
We look at the summary statistics for the OpinPref dataset. Table 8. We see that, interestingly, annotators prefer longer summaries – this is because these summaries contain more specifics and details from the reviews.

### v2

@@ 6 Analysis
We perform a two-fold analysis: (a) First, we see the domain knowledge features influence for $\varphi_{op}$, (b) Second, we see how the ground truth summary and summary from trained models fare on the domain knowledge features. This two-fold analysis helps us understand: (a) which features influence the latent reward model within humans [^9] Note that the trained $\varphi_{op}$ represents latent human reward model. the most, and (b) how the ground truth summary and summary from trained models fare on these influential features. Performing well on influential features would mean the summary aligns well with the latent reward model within humans.
@@ 6.1 Analysis of $\varphi_{op}$
$\varphi_{op}$ model has been trained on a set of features specified by domain experts. We analyze the relative influence of each feature on the final score assigned by $\varphi_{op}$. Doing this helps us understand an approximate importance [^10] We call this approximate importance as the influence of a feature on the output is not necessarily its importance. of each of these features. We do this by varying each feature by $\delta\>(=0.1)$ while keeping the other features constant, over multiple possible values of all features (Equation 6.1).
@@
Figure 6 highlights the features’ relative influence. We see that hallucination is most influential. This aligns with what our human preference annotators report—hallucination in summary is the primary cause of rejection. We see that hallucinations are mostly within the opinions in the summary. This is also reflected in Figure 6: opinion-faithfulness has significant influence. We also see that annotators prefer summaries with more specifics, i.e. they include more aspects: aspect-coverage has significant influence.
@@ 6.2 Analysis of Summaries
We analyze the top-$3$ performing models (in human and Gpt-$4$ evaluations) for the following features: opinion-coverage, opinion-faithfulness, hallucination and relevance. We show the analysis only for the Amazon benchmark in the main manuscript, we include the rest in Appendix J. Table 2 shows the performance on these features. We see that Inductive-Bias model fares much better than the competitors on hallucination (the most influential metric). For relevance, aspect-coverage and opinion-faithfulness, our model is fairly better than the other models.
@@
Models AC $\uparrow$ OPF $\uparrow$ RE $\uparrow$ HL $\uparrow$
IB $\mathbf{3.60}$ $\mathbf{3.93}$ $\mathbf{4.06}$ $\mathbf{4.07}$
SF $3.43$ $3.73$ $4.04$ $3.94$
NM $3.57$ $3.91$ $4.04$ $3.09$
Table 2: Scores on domain knowledge-based features (AC: aspect-coverage, RE: relevance, OPF: opinion-faithfulness, HL: hallucination) on the Amazon benchmark for top-$3$ models (IB: Inductive-Bias, NM: NaiveMean, SF: Synth-Feedback). Note that for hallucination, $\Phi$ gives a higher score for less hallucination in the text.
This shows that our technique helps Inductive-Bias model perform well on features that influence the latent reward model within humans for opinion summarization. This means that our technique helps Inductive-Bias model achieve a significant alignment with the latent reward model. This conclusion verifies our hypothesis (in the domain of opinion summarization): A domain-knowledge infused reward model ($\varphi_{op}$) can help achieve alignment with latent reward model of humans for a task, with modest human preference annotations.
@@ 7 Summary, Conclusion and Future Work
In this work, we propose a novel Reward Modeling technique via Domain Knowledge Infusion. We verify our approach for E-Commerce Opinion Summarization, where we achieve State-of-the-Art, while significantly reducing the amount of human preference annotations required (just $940$ samples). In addition to advancing SOTA and reducing preference annotations, our technique provides another two-fold benefits: (a) No Alignment Tax and (b) Interpretability. Due to the interpretable nature, we find that our model does achieve alignment with human goals for Opinion Summarization through analysis. From the results and analysis, we conclude that Domain Knowledge Infusion into Reward Modeling is a viable solution to reduce human preference annotations for downstream tasks. In the future, we will verify this for other domains.
@@ 8 Ethical Considerations
We contribute two datasets in our work: PromptOpinSumm, OpinPref. These datasets are generated using an open-source model Mistral-$7$B Jiang et al. (2023). We would release the datasets to further research in Opinion Summarization. For the OpinPref, to the best of our knowledge, we have seen that it does not contain any harmful content, such as social biases, stereotypes, etc. However, we have seen that it contains products of explicit nature (sexual products). For the PromptOpinSumm dataset, to the best of our knowledge, there is no presence of harmful content, such as social biases, stereotypes etc. We urge the research community to use the datasets with caution and check for potential harmfulness, based on their use-cases.
@@ 9 Limitations
A limitation of our work is we have tested our approach for one domain: Opinion Summarization. However, we do not believe that this weakens our argument, as we have exhaustively shown that our approach not only advances SOTA but also interpretably achieves alignment with humans. Future work in other domains would help in verifying this claim for other domains. Another limitation is: we see empirically that $\Phi$ works well for Opinion Summarization, to extract the scores for the $7$ features. However, there is no guarantee that such out-of-the-box performance would be reflected in another domain. Some fine-tuning might be necessary.
Appendix A Features for Reward Modeling
We use $7$ domain specific features for the reward model $\varphi_{op}$. We identify these features after extensive discussions with the domain experts. For each feature we prompt Mistral-$7$B to generate a score within 00 and $5$. We give elongated instructions, including rough rubriks, and rules to generate the scores. This is a reason why we use an instruction-tuned model. For each feature, 00 means the model is doing bad on the feature, and $5$ means the model is doing good on the feature. We define all the features below:
@@ Appendix A Features for Reward Modeling
conciseness: This feature considers the conciseness and completeness of the opinion summary. The feature assumes a value $5$ if the summary is concise and complete—not one phrase/sentence can be dropped off. It assumes a value 00 if the summary is totally incomplete, or very verbose.
@@
We also analyze how these features correlate with humans’ judgement of goodness of opinion summaries. We do this by looking at the scores for these features for preferred and dis-preferred summaries in the OpinPref dataset. In Table 3, we see that the preferred summaries clearly have a higher score on all the features, than the dis-preferred ones. This shows that the scores correlate well with humans’ judgement of goodness.
Feature Pref. Dis-pref.
aspect-coverage ($\uparrow$) $3.69$ $2.84$
opinion-faithfulness ($\uparrow$) $4.02$ $3.05$
opinion-coverage ($\uparrow$) $3.92$ $3.22$
conciseness ($\uparrow$) $4.05$ $3.44$
relevance ($\uparrow$) $4.10$ $3.10$
hallucination ($\uparrow$) $3.99$ $2.79$
language-correctness ($\uparrow$) $4.50$ $3.32$
Table 3: Scores for the domain knowledge based features. We see that for all the features, the human preferred (Pref.) summaries have higher scores than the ones rejected by humans (Dis-pref.). This shows that these features correlate well with humans’ judgement of goodness of an opinion summary.
Model-Code Flipkart Flipkart-R Flipkart-Rdq
MedosSiledar et al. (2024) $25.97$ $\mathbf{5.29}$ $\underline{16.05}$ $26.29$ $4.03$ $16.59$ $22.92$ $4.30$ $16.35$
Ours’ DPO $28.85$ $4.10$ $15.55$ $34.23$ $7.86$ $18.62$ $29.96$ $5.25$ $17.28$
Supervised $27.38$ $4.09$ $15.37$ $39.32$ $10.52$ $22.56$ $32.25$ $6.88$ $19.04$
NaiveMean $28.34$ $\underline{4.38}$ $\mathbf{16.20}$ $\mathbf{40.56}$ $10.68$ $22.74$ $\underline{32.57}$ $6.67$ $\underline{19.39}$
Synth-Feedback $26.37$ $4.18$ $15.48$ $38.77$ $\underline{10.99}$ $\underline{22.97}$ $31.04$ $\underline{6.98}$ $18.59$
Inductive-Bias $\mathbf{27.42}$ $4.21$ $15.71$ $\underline{39.10}$ $\mathbf{11.03}$ $\mathbf{23.30}$ $\mathbf{33.08}$ $\mathbf{7.30}$ $\mathbf{19.46}$
Table 4: Reference-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$) for the Flipkart, Flipkart-R and Flipkart-Rdq benchmarks. We see the following things: (a) Our proposed dataset (PromptOpinSumm) leads to marked increased over the SOTA (Medos; by $\sim 6$ R-L points), (b) Inductive-Biasproves to be the winner in all the four scenarios: Scene-I, Scene-II, Scene-III and Scene-IV (Section 5.1), proving the efficacy of our technique. We also see that for Flipkart benchmark, despite the shortcomings, our models perform similar to the SOTA.
Model-Code Oposum+ Oposum-R Oposum-Rdq
Prior Works MeanSum Chu and Liu (2018) $26.25$ $4.62$ $16.49$- -- -- -- -- --
CopyCat Bražinskas et al. (2020) $27.98$ $5.79$ $17.07$ $22.41$ $2.30$ $13.94$ $22.38$ $2.03$ $14.06$
PlanSum Amplayo and Lapata (2020) $30.26$ $5.29$ $17.48$ $22.37$ $2.05$ $13.32$ $22.64$ $2.25$ $13.71$
MultimodalSum Im et al. (2021) $33.08$ $7.46$ $19.75$ $23.35$ $2.98$ $14.53$ $23.73$ $2.80$ $14.70$
Op-Sum-GenSiledar et al. (2023a) $36.44$ $8.50$ $22.03$ $25.65$ $3.56$ $15.83$ $24.66$ $3.25$ $15.54$
MedosSiledar et al. (2024) $\mathbf{36.57}$ $\underline{8.79}$ $\mathbf{21.35}$ $26.82$ $3.67$ $15.92$ $26.32$ $3.34$ $16.10$
Ours’ DPO $27.64$ $7.34$ $16.50$ $33.69$ $6.62$ $18.55$ $30.95$ $5.89$ $17.60$
Supervised $30.57$ $8.02$ $16.90$ $38.32$ $9.10$ $20.35$ $35.69$ $8.17$ $19.28$
NaiveMean $31.47$ $8.0$ $16.99$ $40.16$ $9.84$ $21.74$ $35.90$ $8.33$ $20.13$
Synth-Feedback $\underline{31.66}$ $\mathbf{8.86}$ $\underline{17.91}$ $\underline{41.32}$ $\mathbf{10.40}$ $\mathbf{22.23}$ $\mathbf{37.85}$ $\underline{8.94}$ $\underline{20.71}$
Inductive-Bias $31.15$ $8.15$ $17.46$ $\mathbf{41.58}$ $\underline{10.32}$ $\underline{22.02}$ $\underline{37.56}$ $\mathbf{9.21}$ $\mathbf{20.88}$
Table 5: Reference-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$) for the Oposum+, Oposum-R and Oposum-Rdq benchmarks. We see the following things: (a) Our proposed dataset (PromptOpinSumm) leads to marked increased over the SOTA (Medos; by $\sim 6$ R-L points), (b) Inductive-Biasproves to be the winner in almost all of the four scenarios: Scene-I, Scene-II, Scene-III and Scene-IV (Section 5.1), proving the efficacy of our technique. We also see that for Oposum+ benchmark, our models lag behind. However, this is expected, as we highlight in Section 3.2.
Model Code Amazon Amazon-R Amazon-Rdq Oposum+ Oposum-R Oposum-Rdq Flipkart Flipkart-R Flipkart-Rdq
Op-Sum-Gen $\mathbf{88.78}$ $86.94$ $86.76$ $\mathbf{86.63}$ $86.96$ $86.95$- -- --
Siledar et al. (2023a)
DPO $86.45$ $86.60^{\dagger}$ $86.37^{\dagger}$ $84.39$ $87.35^{*}$ $86.90$ $83.75$ $86.61$ $85.40$
Supervised $87.79$ $88.23^{*}$ $87.76^{*}$ $85.13$ $88.59^{*}$ $88.02^{*}$ $84.21$ $88.11$ $86.40$
NaiveMean $87.95$ $\underline{88.29}^{*}$ $\underline{87.81}^{*}$ $85.25$ $88.96^{*}$ $88.39^{*}$ $\underline{84.32}$ $\underline{88.29}$ $\underline{86.52}$
Synth-Feedback $87.81$ $88.28^{*}$ $87.74^{*}$ $85.22$ $\underline{89.08}^{*}$ $\underline{88.45}^{*}$ $84.27$ $88.28$ $86.49$
Inductive-Bias $\underline{87.98}$ $\mathbf{88.41}^{*}$ $\mathbf{88.16}^{*}$ $\underline{85.33}$ $\mathbf{89.09}^{*}$ $\mathbf{88.46}^{*}$ $\mathbf{84.33}$ $\mathbf{88.34}$ $\mathbf{86.61}$
Table 6: BertScore evaluation results on the $9$ benchmark datasets. We observe a similar trend as Rouge evaluations: SOTA is better than our models for the Amazon and Oposum+ benchmarks, which is expected (Section 3.2). For the rest of the datasets, we see that our models are significantly better. We do not include SOTA results for Flipkart-based benchmarks, as Op-Sum-Gen only provide aspect-specific summaries for the same. denotes gain is statistically significant compared to SOTA with significance level $1\%$, denotes gain is statistically significant compared to SOTA with significance level $5\%$.
Appendix B RLHF Training Pipeline
Using the trained reward model, we follow a similar training pipeline as Bai et al. (2022a); Ouyang et al. (2022), with a modification: Limited Trajectory Reinforcement Learning. Computing the transformation $\Phi$ for each generation online (during training) is expensive, especially with limited compute resources. To circumvent this, we limit the trajectories that are explored by our policy, $\pi_{\theta}$. Specifically, we limit it to the Good-Sum, SBad-Sum and VBad-Sum trajectories in the PromptOpinSumm dataset. Having varying levels of quality in PromptOpinSumm is of use here—it lets the model still explore trajectories of several quality. Thus, we have an offline experience buffer, with $\Phi$ precomputed, for $\pi_{\theta}$ learn from.
We use Proximal Policy Optimization (PPO) Schulman et al. (2017) to train our model (Equation B). For each training step, we sample ($R,s,\Phi(R,s)$) tuples from PromptOpinSumm. We use the trained $\varphi_{op}$ to compute the reward for $s$ ($=\varphi_{op}\big{(}\Phi(R,s)\big{)}$). PPO uses this to update the log probability assigned by $\pi_{\theta}$. We parameterize $\pi_{\theta}$ using a Transformer model, which takes reviews as input, and generates an opinion summary.
Appendix C Additional Automatic Evaluation Results
In addition to the Amazon-based benchmarks (Table 1), we also report results for Flipkart and Oposum+ based benchmarks (Tables 4 and 5). As before, we see that Inductive-Bias is almost always the winner. As before, we draw similar conclusions for Scene-I, Scene-II and Scene-III: Inductive-Bias wins, further strengthening the conclusion that our methodology is effective. We also see that, inspite of the shortcomings of the Flipkart benchmark, our models perform similar to the SOTA.
We also include BertScore evaluations for all the $9$ benchmarks in Table 6. We see similar trends as Rouge Evaluation: our models are significantly better than the SOTA in majority of the benchmarks.
For a qualitative understanding, we include generations from several models on a randomly picked sample from the Amazon benchmark in Table 14.
Appendix D Details on the Benchmark Datasets
In this section we discuss details about the benchmarks, such as the domain of the products, summary statistics and finally highlight some shortcomings in the Amazon, Oposum+ and Flipkart datasets. Amazon has reviews for $32$ products from $4$ domains: “electronics”, “home & kitchen”, “personal care”, and “clothing, shoes & jewellery”. Oposum+ has reviews for $60$ products from $6$ domains: “laptop bags”, “bluetooth headsets”, “boots”, “keyboards”, “television”, and “vacuums”. Flipkart has reviews for $147$ products from $3$ domains: “laptops”, “mobiles”, and “tablets”. Table 7 includes summary statistics for the benchmarks.
Characteristic Oposum+ Amazon Flipkart
# domains $6$ $4$ $3$
# products $60$ $32$ $147$
# reviews $10$ $8$ $10$
per product
# summaries $3$ $3$ $1$
per product
Table 7: Statistics of the benchmark datasets. Oposum+ represents the statistics of all Oposum+ based benchmarks (Oposum+, Oposum-R and Oposum-Rdq). Similar is the case for Amazon and Flipkart.
Finally, now we highlight the shortcomings of the benchmark datasets in the rest of the discussion.
@@ Appendix D Details on the Benchmark Datasets
Amazon: Bražinskas et al. (2020) designed the test-set in such a way that the summary has to read like a review, for instance, summary would contain ‘I think the quality has come down over the years.’, instead of ‘Users think that quality has come down over years’. Due to this writing style, the summaries read like reviews and are often in first person—high overlap would not necessarily mean a better summary, it would rather mean a better review.
@@
Appendix E Implementation Details
@@
Appendix F Generated Summary Lengths
We analyze the generation lengths of the models, and the ground truth summary. Table 10 lists the summary lengths. We see that theDPO model generates very verbose summary. Additionally, we also see that the Inductive-Bias model generates very concise summaries.
Model Amazon Oposum+ Flipkart
@@ Appendix F Generated Summary Lengths
NaiveMean $91.09$ $114.67$ $75.48$
Synth-Feedback $80.31$ $115.37$ $71.11$
@@
Inductive-Bias $81.62$ $88.63$ $73.57$
Supervised $81.31$ $117.03$ $74.56$
DPO $138.50$ $141.50$ $131.40$
Table 10: Generation Length Statistics: number of words in summaries. We use Nltk to tokenize the text.
@@
Table 11: Example summaries from PromptOpinSumm.
Appendix G Details on PromptOpinSumm
Here we provide more details on the generated PromptOpinSumm dataset. Table 13 includes summary statistics of the generated dataset. We include an example from PromptOpinSumm dataset (Table 11). We show one sample from Good-Sum, SBad-Sum and VBad-Sum quality each. We do not include the reviews for conciseness. However, we incluce the salient aspects of the reviews. The reviews talk about the following things:
@@ Appendix G Details on PromptOpinSumm
Good-Sum manages to discuss all of these things—it was able to detect that the cover does not fit specific models (highlighted in blue). It also detects that a few users like the “personal touch addition” factor of the cover.
VBad-Sum is also a good representation of a bad summary—it totally ignores the “positive” aspects of the product and presents the “negative” aspects only. It fails at Aspect Coverage, Opinion Faithfulness and Opinion Coverage.
@@
Appendix H Statistics of the OpinPref dataset
We look at the summary statistics for the OpinPref dataset. Table 12. We see that, interestingly, annotators prefer longer summaries—this is because these summaries contain more specifics and details from the reviews.